You deposit $1230 into a savings account earning 6.5% interest, compounded semi-annually. How many periods does it take to accumulate amount of $3,109.69 in the account?
Answer by stanbon(75887)

You deposit $1230 into a savings account earning 6.5% interest, compounded semi-annually. How many periods does it take to accumulate amount of $3,109.69 in the account?
------------
A(t) = Ao(1+(r/n))^(nt)
3109.69 = 1230(1+(0.065/2))^(2t)
----
2.5282 = 1.0325^(2t)
---
(2t)log(1.0325) = log(2.5282)
2t = 29
---
t = 14.5 years
=====================
Cheers,
Stan H.